### v1

# New Types of Hydrogenlike matter Composed of Electron(s) and Meson(s)

Abstract
In the present work, we predict the existence of new types of hydrogenlike matter, including hydrogenlike atoms ($\pi^{+}e^{-}$, $K^{+}e^{-}$, $D^{+}e^{-}$), hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$) and hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$). By solving the Schrödinger equation, the binding energy of hydrogenlike atoms is obtained as $E_{n}=-\frac{1}{2n^{2}}$. For hydrogenlike molecular ions and molecules, the variational method is employed to calculate the binding energies, i.e., $E_{+}=-0.587$ and $E_{0}=-1.139$ for hydrogenlike molecular ions and molecules, respectively. And the bond lengths for hydrogenlike molecular ions and molecules are also calculated, whose values are $2.003$ and $1.414$, respectively. Here all the quantities are in atomic unit for convenience. In additon, the strong interaction between the two constituent mesons is considered in our calculations, where we find that its influence on the hydrogenlike molecular ions and molecules can be neglected. Hopefully, these types of matter would be observed in the future with the improvement of accuracy in the high energy physical experiments.
I Introduction
The exploration of microscopic matter and their inner structure can increase our understanding of nature, which is always full of interests and challenges.
J. Dalton from England proposed the atomic theory in 1803, and opened a new era in modern chemistry. Dalton’s theory not only explained many chemical phenomena and chemical laws successfully, but also revealed their inner connections, transforming ancient naive atomic ideas into a scientific atomic theory. Soon after that, J. L. Gay-Lussac from France summarized experimental results of various gas reactions and drew an important conclusion: when gases react with each other, they often combine in simple ratios, and the change in volume of the combined gas also has a simple proportional relationship with the volumes of the reacting gases. Lussac’s conclusion leaded to a problem of half an atom, which contradicts the indivisibility of atoms in atomic theory and is also impossible in chemical reactions. In order to solve this problem, Amedeo Avogadro from Italy examined Lussac’s experiments and his academic dispute with Dalton, and proposed a kind of particle of a new divisible level-molecules. In nearly half a century, Avogadro’s molecular theory was not recognized by the academic community. As the rapidly development of modern inorganic and organic chemistry during this period, Italian scientist Stanislao Cannizzaro provided theoretical justification for the concept of molecules, organically integrating Dalton’s atomic theory with Avogadro’s molecular theory, and systematically proposing the scientific atomic-molecular theory, which was ultimately established in academic community.
At the turn of 19th and 20th centuries, the discovery of X-rays, radioactivity, along with the development of spectroscopy, the proposal of empirical formulas for spectral laws, and the electron-theoretical explanation of the Zeeman effect had changed the traditional concepts that atoms were indivisible and elements unchangeable. Scientists successively proposed several models of atomic structure, but they were all successively disproved by experiments. In 1911, Ernest Rutherford designed the $\alpha$ particles scattering experiment, and revealed that most $\alpha$ particles passed directly through the gold foil, while a small fraction were deflected at large angles of even almost $180^{\circ}$ Rutherford (1911). In order to explain this phenomenon, Rutherford proposed the nuclear structure model, whose core idea is that there is an extremely small atomic nucleus at the center of an atom, and all electrons revolve around the atomic nucleus. However, it was impossible to understand the stability of atoms, and to explain why excited atoms emit discrete spectral lines rather than continuous ones. So in 1913, Bohr proposed a new model of atomic structure, which is based on stationary state hypothesis, frequency rule, and quantization condition of angular momentum Bohr (1913a, b, c). In 1924, inspired by the wave-particle duality of light, Louis de Broglie proposed that physical particles such as electrons also exhibit wave-particle duality de Broglie (1925). This viewpoint directly leaded to the birth of quantum mechanics, which provides a theoretical foundation for understanding the structure of atoms.
About 20 years after the nuclear structure model was proposed by Rutherford, a new kind of neutral particle named neutron was observed by J. Chadwick Chadwick (1932), which was a milestone in the history of nuclear physics development. Then Heisenberg and Iwanenko proposed that nuclear is make of protons and neutrons Heisenberg (1932); Ivanenko (1932). After that, nuclear physics, as an independent research field, continued to develop in experiments, theories and applications.
However, another puzzle was how protons and neutrons could be tightly bound together in a nucleus with a size of $10^{-13}$ centimeters. In 1935, Hideki Yukawa proposed that there exists a strong binding force between protons and neutrons within the atomic nucleus by exchanging an unknown meson Yukawa (1935). This meson, called pion ($\pi$), has a mass of approximately $100-200$ MeV. It was not until 1947 that C. Powell et al. discovered the $\pi$ meson Lattes et al. (1947a, b). In the 1950s, with the development of high-energy accelerators, a large number of particles directly involved in strong interactions were discovered in accelerator experiments, and they had extremely short lifetimes. By the early 1960s, more than 100 “elementary” particles had been discovered in nature, which could be classified into two categories according to their interactions:
  • Hadrons: Particles directly participating in strong interactions, such as protons, neutrons, $\pi$ mesons, strange particles and a series of resonant particles.
  • Leptons: Particles do not directly involved in strong interactions but only in electromagnetic and weak interactions, such as electrons, muons and neutrinos.
High-energy physics experiments further revealed that hundreds of hadrons were not “fundamental” but had internal structures. In 1964, M. Gell-Mann and Zweig proposed the quark model Gell-Mann (1964); Zweig (1964a, b), regarding quarks as a new hierarchial level in the structure of matter. All the hadrons known at that time were composed of three fundamental quarks: the up quark (u), down quark (d), and strange quark (s). Numerous high-energy physics experiments in the 1960s confirmed the existence of quarks and the success of the quark model. Subsequent discoveries expanded the quark family. In 1974, Sounuel C. C. Ting and Richter discovered the charm quark (c) Aubert et al. (1974); Augustin et al. (1974). In 1977, Lederman et al. discovered the bottom quark (b) Innes et al. (1977). In 1995, the top quark (t) was discovered Abe et al. (1994a, b). The six quarks together with their corresponding antiquarks serve as the “building blocks” for all hundreds of hadrons. Meanwhile, the discovery of leptons also reached six types: electrons, electron neutrinos, muon, muon neutrinos, tau lepton, and tau neutrinos. Thus, quarks and leptons are currently recognized as the deepest and smallest constituent components of matter at our present level of understanding. With the further development of accelerator and detection technologies, nowadays many new particles have been observed including Higgs, the traditional mesons and baryons, tetraquark states ($Z_{c}(3900)$, $T_{cc}^{+}$, $\cdots$), and pentaquark states ($P_{c}(4312)$, $P_{c}(4380)$, $P_{c}(4440)$, $P_{c}(4457)$, $\cdots$), see Refs. Chen et al. (2016); Liu et al. (2019); Olsen (2015); Hosaka et al. (2016); Lebed et al. (2017); Esposito et al. (2017); Guo et al. (2018); Ali et al. (2017); Karliner et al. (2018); Albuquerque et al. (2019); Brambilla et al. (2020); Guo et al. (2020); Meng et al. (2023); Bicudo (2023); Liu et al. (2024); Chen et al. (2023); Dong et al. (2021, 2017); Yang et al. (2020); Drenska et al. (2010); Briere et al. (2016); Richard (2016); Mezzadri and Spataro (2022); Johnson et al. (2024) for reviews.
Besides, some kinds of composite particles have also been studied, which involve few-lepton systems (see Ref. Ma et al. (2025) and the references there in), muonic atoms, hadronic atoms (see the reviews Friedman and Gal (2007); Batty et al. (1997); Gal et al. (2016)), and $\pi-\mu$ Coulomb bound states Coombes et al. (1976). In Refs. Abdel-Raouf et al. (1998); Rebane (2012a); Bressanini et al. (1998); Rebane (2012b); Richard (1994); Varga et al. (1997); Mitroy and Novikov (2004); Rebane (2012c); Bressanini et al. (1997); Rebane and Markovskiĭ (2001); Rebane and Markovski (2003); Gridnev and Greiner (2005); Van Hooydonk (2005); El-Khateeb (2007); Suzuki et al. (1998); Richard (2002); Armour et al. (2005); Emami-Razavi (2015); Emami-Razavi and Darewych (2021), four-body-exotic systems were discussed, including the predictions of $\pi^{+}\pi^{+}e^{-}e^{-}$ and $K^{+}K^{+}e^{-}e^{-}$ bound states Rebane (2012a, b). Hydrogenlike charm-pion or charm-kaon matter was predicted in Ref. Luo et al. (2023).
[…]
This paper is organized as follows. After the introduction, the theoretical framework is described in Sec. II, which contains the calculations of $Me^{-}$, $MMe^{-}$ and $MMe^{-}e^{-}$ systems ($M$ = $\pi^{+}$, $K^{+}$, $D^{+}$). After that, we show the numerical results in Sec. III and give a discussion in Sec. IV whether the strong interaction affect the systems. In Sec. V, a short summary is given.
II FORMALISM
In this work, we will consider three new types of matter composed of meson(s) ($\pi^{+}/K^{+}/D^{+}$) and electron(s). These constituent masses are similar to those of the hydrogen atom, molecular ion and molecule. In order to derive the binding energy, we use the method by analogy with the hydrogen cases, which are shown in the following.
II.1 Hydrogenlike atoms ($\pi^{+}e^{-},K^{+}e^{-}$, $D^{+}e^{-}$)
We present the calculation procedure for the hydrogenlike atom. Adopting atomic units (a.u.) for this discussion, the Hamiltonian of the hydrogenlike atom, i.e., $\pi^{+}e^{-},K^{+}e^{-}$ and $D^{+}e^{-}$ reads
$H$ $=$ $-\frac{1}{2}\nabla^{2}-\frac{1}{r}$ (1)
$=$ $-\frac{1}{2}\frac{1}{r^{2}}~{}\frac{\partial}{\partial r}~{}r^{2}~{}\frac{\partial}{\partial r}+\frac{1}{2}~{}\frac{1}{r^{2}}~{}\hat{\vec{L}}^{2}-\frac{1}{r},$
where $\hat{\vec{L}}$ is the angular momentum operator, i.e.,
$\hat{\vec{L}}$ $=$ $\hat{\vec{r}}\times\hat{\vec{p}}=-i\vec{r}\times\nabla$ (2)
$=$ $-i~{}\bigg{(}\vec{e_{\varphi}}~{}\frac{\partial}{\partial\theta}-\vec{e_{\theta}}~{}\frac{1}{\sin{\theta}}~{}\frac{\partial}{\partial\varphi}\bigg{)}.$
and
$\hat{\vec{L}}^{2}$ $=$ $-\bigg{(}\frac{1}{\sin{\theta}}~{}\frac{\partial}{\partial\theta}~{}\sin{\theta}~{}\frac{\partial}{\partial\theta}+\frac{1}{\sin^{2}{\theta}}~{}\frac{\partial^{2}}{\partial\varphi^{2}}\bigg{)}.$ (3)
Note here that in a.u., $m_{e}=e=\hbar=1$. After solving the stationary-state Schrödinger equation, we drive
$E_{n}$ $=$ $-\frac{1}{2n^{2}},$ (4)
$\psi_{nlm}(r,\theta,\varphi)$ $=$ $R_{nl}(r)~{}Y_{lm}(\theta,\varphi),$ (5)
where the radial wave function is
$R_{nl}(r)$ $=$ $\frac{2}{n^{2}(2l+1)!}\bigg{[}\frac{(n+l)!}{(n-l-1)!}\bigg{]}^{1/2}\xi^{l}$ (6)
×$\times F(-n_{r},2l+2,\xi)e^{-\xi/2},$
the Spherical Harmonics is $Y_{lm}(\theta,\varphi)$, and
$F(a,b,\chi)$ $=$ $1+\frac{a}{b}\chi+\frac{a(a+1)}{b(b+1)}\frac{\chi^{2}}{2!}+\frac{a(a+1)(a+2)}{b(b+1)(b+2)}\frac{\xi^{3}}{3!}$ (7)
+$\cdots$+ ⋯
is the Confluent hypergeometric function, $n=n_{r}+l+1=1,2,3,\cdots$ is the principal quantum number, $\xi=\frac{2r}{n}$, $l=0,1,2,\cdots$ is the angular momentum quantum number, $m=0,\pm 1,\pm 2,\cdots$ is the magnetic quantum number. Specially, the energy and wave function of the ground eigenstate are
$E_{1}$ $=$ $-\frac{1}{2},$ (8)
$\psi_{100}(r,\theta,\varphi)$ $=$ $\bigg{(}\frac{1}{\pi}\bigg{)}^{1/2}e^{-r}.$ (9)
II.2 Hydrogenlike molecular ions ($\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$, $D^{+}D^{+}e^{-}$)
According to the Born-Oppenheimer approximation, the relative distance between the two mesons can be viewed as parameter, not the dynamical variable, such that the Hamiltonian of the systems can be expressed as
$H$ $=$ $-\frac{1}{2}\nabla^{2}-\frac{1}{r_{a}}-\frac{1}{r_{b}}+\frac{1}{R}$ (10)
$=$ $H_{e}+\frac{1}{R},$
where the spin of the electron is neglected, $r_{a}$ and $r_{b}$ are the relative distances between the electron and the two mesons, respectively, $H_{e}\equiv-\frac{1}{2}\nabla^{2}-\frac{1}{r_{a}}-\frac{1}{r_{b}}$ depicts the motion of the electron in the two-center potential.
Like the case of hydrogen molecular ion, we adopt the variational method to calculate the binding energy of the systems. In this work, we only focus on the ground eigenstate case. In light of that the electron is only affected by the attractive Coulomb force of the two identical mesons, it is naturally to choose the tried wave function as the superposition as follows
$\psi=C_{a}\psi_{a}+C_{b}\psi_{b},$ (11)
where
$\psi_{a}=\bigg{(}\frac{\lambda^{3}}{\pi}\bigg{)}^{1/2}e^{-\lambda r_{a}},$ (12)
$\psi_{b}=\bigg{(}\frac{\lambda^{3}}{\pi}\bigg{)}^{1/2}e^{-\lambda r_{b}},$ (13)
with $\lambda$ is the variational parameter. If $\lambda$ is fixed as $\lambda=1$, $\psi_{a}$ and $\psi_{b}$ become the ground-state wave function of hydrogenlike atom, i.e., Eq. (9).
Since the potential of the system is invariant when performing the reflection with respect to the midpoint of the two mesons, the state of the electron can be classified according to the reflection symmetry. Here we have $C_{a}=C_{b}\equiv C_{+}$ for even parity, and $C_{a}=-C_{b}\equiv C_{-}$ for odd parity. So the tried wave function can be expressed as
$\psi_{\pm}=C_{\pm}(\psi_{a}\pm\psi_{b}).$ (14)
From the normalization condition, we obtain
$C_{\pm}=\frac{1}{\sqrt{2}}\bigg{\{}1\pm\big{[}\frac{1}{3}(\lambda R)^{2}+\lambda R+1\big{]}e^{-\lambda R}\bigg{\}}^{-1/2}.$ (15)
Using Eq. (14), we get the expression for the average value of the energy
$E_{\pm}$ $=$ $\langle\psi_{\pm}\mid H\mid\psi_{\pm}\rangle$
$=$ $\langle\psi_{\pm}\mid H_{e}\mid\psi_{\pm}\rangle+\frac{1}{R}$
$=$ $2C^{2}_{\pm}\bigg{\{}\frac{\lambda^{2}}{2}-\lambda-\frac{1}{R}+\frac{1}{R}(\lambda R+1)e^{-2\lambda R}$
±$\pm\big{[}\frac{\lambda^{3}R}{2}+\frac{\lambda^{2}}{2}-2\lambda^{2}R-2\lambda-\frac{\lambda^{4}R^{2}}{6}\big{]}e^{-\lambda R}\bigg{\}}+\frac{1}{R}.$
The variational parameters $\lambda$ and $R$ are determined by the variational extremum condition, i.e.,
$\frac{\partial E_{\pm}}{\partial R}=0,~{}~{}\frac{\partial E_{\pm}}{\partial\lambda}=0.$ (17)
II.3 Hydrogenlike molecules ($\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$, $D^{+}D^{+}e^{-}e^{-}$)
In 1927, Heitler and London successfully applied quantum mechanics to study the structure of hydrogen molecule Heitler and London (1927), and pioneered the field of quantum chemistry.
Similar to the hydrogen molecule system, the hydrogenlike molecule system consists of two identical mesons and two electrons. According to Born-Oppenheimer approximation, the motion of the two mesons can be neglected, such that the hydrogenlike molecule is simplified as two-electron problem, of which the Hamiltonian reads
$H$ $=$ $-\frac{1}{2}(\nabla_{1}^{2}+\nabla_{2}^{2})-\frac{1}{r_{a1}}-\frac{1}{r_{a2}}-\frac{1}{r_{b1}}-\frac{1}{r_{b2}}$ (18)
$+\frac{1}{r_{12}}+\frac{1}{R}.$
The distance $R$ between the two mesons is treated as a parameter in our calculation. And next we utilize the variational method to get the eigenenergy and wave function of the ground state. In view that when $R$ becomes very large, the ground-state wave function of the hydrogenlike molecule can approximately written as the product of the ground-state wave functions of the two hydrogenlike atoms, considering the exchange symmetry of the two electrons, the tried wave function can be chosen as
$\psi_{+}(1,2)$ $=$ $\phi_{0}(1,2)~{}\chi_{0}(S_{1z},S_{2z}),$ (19)
$\psi_{-}(1,2)$ $=$ $\phi_{1}(1,2)~{}\chi_{1}(S_{1z},S_{2z}),$ (20)
$\phi_{0}(1,2)$ $=$ $C_{S}\bigg{[}\psi(r_{1a})\psi(r_{2b})+\psi(r_{2a})\psi(r_{1b})\bigg{]},$ (21)
$\phi_{1}(1,2)$ $=$ $C_{A}\bigg{[}\psi(r_{1a})\psi(r_{2b})-\psi(r_{2a})\psi(r_{1b})\bigg{]},$ (22)
where $\chi_{0}(S_{1z},S_{2z})$ and $\chi_{1}(S_{1z},S_{2z})$ are the spin wave functions of singlet and triplet, respectively. Since the Hamiltonian $H$ does not depend on spin, we do not consider spin wave function in the calculation of variational method. Taking into account the influence on the electron of the mesons and another electron, the single-electron wave functions can be chosen as
$\psi(r)=\bigg{(}\frac{\lambda^{3}}{\pi}\bigg{)}^{1/2}e^{-\lambda r},$ (23)
where $\lambda$ is the variational parameter. $C_{S}$ and $C_{A}$ are determined by the normalization conditions, i.e.,
$C_{S}$ $=$ $\frac{1}{\sqrt{2}}\bigg{[}1+\bigg{(}\frac{1}{3}\lambda^{2}R^{2}+\lambda R+1\bigg{)}^{2}e^{-2\lambda R}\bigg{]}^{-1/2},$ (24)
$C_{A}$ $=$ $\frac{1}{\sqrt{2}}\bigg{[}1-\bigg{(}\frac{1}{3}\lambda^{2}R^{2}+\lambda R+1\bigg{)}^{2}e^{-2\lambda R}\bigg{]}^{-1/2}.$ (25)
The average value of the energy are obtained as
$E_{0}$ $=$ $\langle\phi_{0}(1,2)\mid H\mid\phi_{0}(1,2)\rangle$ (26)
$=$ $4C_{s}^{2}I_{H1}+4C_{s}^{2}I_{e1}+4C_{s}^{2}I_{H2}S_{1}+4C_{s}^{2}I_{e2}S_{1}$
$+2C_{s}^{2}I_{DCC}+2C_{s}^{2}I_{DCE}+\frac{1}{R},$
$E_{1}$ $=$ $\langle\phi_{1}(1,2)\mid H\mid\phi_{1}(1,2)\rangle$ (27)
$=$ $4C_{A}^{2}I_{H1}+4C_{A}^{2}I_{e1}-4C_{A}^{2}I_{H2}S_{1}-4C_{A}^{2}I_{e2}S_{1}$
$+2C_{A}^{2}I_{DCC}-2C_{A}^{2}I_{DCE}+\frac{1}{R},$
where
$I_{DCE}$ $=$ $\frac{6}{5R}\bigg{\{}-S_{2}^{2}E_{i}(4\lambda R)+2S_{1}S_{2}E_{i}(2\lambda R)+S_{1}^{2}(\gamma+\ln{(\lambda R)})\bigg{\}}$ (28)
$+\frac{e^{-2\lambda R}}{R}\bigg{[}-\frac{1}{15}(\lambda R)^{4}-\frac{3}{5}(\lambda R)^{3}-\frac{23}{20}(\lambda R)^{2}+\frac{5}{8}\lambda R\bigg{]},$
$I_{DCC}$ $=$ $-\lambda\bigg{(}\frac{\lambda^{2}R^{2}}{6}+\frac{3}{4}\lambda R+\frac{11}{8}+\frac{1}{\lambda R}\bigg{)}e^{-2\lambda R}+\frac{1}{R},$ (29)
$I_{H1}$ $=$ $\frac{\lambda^{2}}{2}-\lambda,$ (30)
$I_{H2}$ $=$ $\bigg{[}\frac{\lambda^{3}R}{2}+\frac{\lambda^{2}}{2}-\lambda^{2}R-\lambda-\frac{\lambda^{4}R^{2}}{6}\bigg{]}e^{-\lambda R},$ (31)
$I_{e1}$ $=$ $-\frac{1}{R}\bigg{[}-(\lambda R+1)e^{-2\lambda R}+1\bigg{]},$ (32)
$I_{e2}$ $=$ $-\lambda(\lambda R+1)e^{-\lambda R},$ (33)
$S_{1}$ $=$ $e^{-\lambda R}\bigg{(}\frac{1}{3}\lambda^{2}R^{2}+\lambda R+1\bigg{)},$ (34)
$S_{2}$ $=$ $e^{\lambda R}\bigg{(}\frac{1}{3}\lambda^{2}R^{2}-\lambda R+1\bigg{)},$ (35)
[…]
The parameters $\lambda$ and $R$ are also determined by the variational extremum condition,
$\frac{\partial E_{0,1}}{\partial R}=0,~{}~{}\frac{\partial E_{0,1}}{\partial\lambda}=0.$ (37)
III NUMERICAL RESULTS
[…]
$E_{1}=-\frac{1}{2},E_{2}=-\frac{1}{8},E_{3}=-\frac{1}{18},\cdots$ (38)
which is the same as that of hydrogen atom, when neglecting the motion of the meson or the proton. Beside the binding energy of $Me^{-}$ systems, we also plot the radial probability density as well as the electron cloud of $n=1,2,3$ and $l=0$ state, which is shown in FIG. 1 and FIG. 2 (a), (b), (c). Due to the electromagnetic interaction between $M$ and $e^{-}$ is the same as that of proton and $e^{-}$ in hydrogen atom, the wave function of electron in $Me^{-}$ systems is consistent with that of in hydrogen atom.
Figure 1: The Radial probability density of the 1S, 2S and 3S state with $\chi_{n0}=rR_{n0}(r)$.
Figure 2: (color online) The electron cloud of hydrogenlike atoms and molecular ions in $xy$ plane. (a), (b), (c) and (d) corresponds to $Me^{-}\ (1S)$, $Me^{-}\ (2S)$, $Me^{-}\ (3S)$ and $MMe^{-}e^{-}$ with $M=\pi^{+},K^{+},D^{+}$, respectively.
For $MMe^{-}$ system of parity odd state, the average value of energy $E_{-}$ does not have minimum value. If $\lambda$ is fixed, we clearly see in FIG. 3 that $E_{-}$ is monotonically decreasing when the distance between the two mesons is increasing. In this case, the electron and two mesons can not form a bound state.
For $MMe^{-}$ system of even parity, there is a minimum value of $E_{+}$, i.e., $E_{+}=-0.587$, where the variational parameters $\lambda=1.238$ and $R=2.003$. So the two identical mesons and the electron can form a bound state. This can be easily understood: since $\psi_{+}$ is symmetric when exchanging the two mesons, there is a large distribution of electron near the midpoint of the mesons, forming a cloud with negative charge (see FIG. 2 (d)); and the electron cloud has an attractive interaction on the mesons with positive charge.
For the hydrogenlike molecule, there is no minimum value for $E_{1}$ which can be seen in FIG. 3. However, $E_{0}$ has a minimum $E_{0}=-1.139$ at $\lambda=1.166$, $R=1.414$. In the limit $\lambda\to 1$ and $R\to\infty$, we obtain from Eq. (26) that $E_{0}=-1$. The picture is that a hydrogen molecule becomes two independent hydrogen atoms in this limit, and the energy of the electrons is the sum of the binding energies ($E_{e}=-\frac{1}{2}$ for each atom) of the atoms. Despite the fact that a hydrogenlike molecule is make of two neutral hydrogenlike atom, they can form a bound state.
Figure 3: The line shapes of the $E_{+}$, $E_{-}$, $E_{0}$ and $E_{1}$. The variational parameter $\lambda$ is fixed as 1.238 for $E_{+}$, $E_{-}$, and as 1.166 for $E_{0}$, $E_{1}$.
For hydrogenlike molecule, the spin wave function of the electron pair is antisymmetric, i.e., the two electrons are in a spin singlet, whereas the spatial wave function is symmetric. So the electrons can yet close to each other, which means they can form electron cloud with relatively higher density. And the electron cloud has a strong attractive force on the two mesons, such that the two hydrogenlike atoms combine together. The structure that the electron pair is shared by both mesons and their spins are in a singlet is so called covalent boud.
IV Discussion: Does the strong interaction affect the systems?
In this section, we discuss whether or not the strong interaction between the mesons impact on the systems considered in our work.
The corresponding Lagrangian are constructed by the hidden local symmetry and chiral symmetry, in which the terms needed in the present work are show below,
$\mathcal{L}_{V\Phi\Phi}$ $=$ $-\frac{i}{\sqrt{2}}\langle[\Phi,\partial_{\mu}\Phi]V^{\mu}\rangle,$ (39)
$\mathcal{L}_{VPP}$ $=$ $ia_{1}(PV_{\mu}D^{\mu}P^{\dagger}-D^{\mu}PV_{\mu}P^{\dagger}).$ (40)
The matrices of $\Phi$, $V_{\mu}$ and $P$ read
$\Phi$ $=$ $\sqrt{2}\left(\begin{array}[]{ccc}\frac{\sqrt{3}\pi^{0}+\eta+\sqrt{2}\eta^{\prime}}{\sqrt{6}}&\pi^{+}&K^{+}\\ \pi^{-}&\frac{-\sqrt{3}\pi^{0}+\eta+\sqrt{2}\eta^{\prime}}{\sqrt{6}}&K^{0}\\ K^{-}&\bar{K}^{0}&\frac{-2\eta+\sqrt{2}\eta^{\prime}}{\sqrt{6}}\end{array}\right),$ (44)
$V_{\mu}$ $=$ $\frac{g_{V}}{\sqrt{2}}\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}(\rho^{0}+\omega)&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{1}{\sqrt{2}}(\rho^{0}-\omega)&K^{*0}\\ K^{*-}&\bar{K}^{*0}&\phi\end{array}\right)_{\mu},$ (48)
$P$ $=$ $\left(D^{0},D^{+},D_{s}^{+}\right).$ (49)
The constant $g_{V}$ is determined from the experimental value of $g_{\rho\pi\pi}$, i.e., $g_{V}=5.80\pm 0.9$ Harada and Yamawaki (2003). Comparing $\mathcal{L}_{VPP}$ with the Lagrangian in Ref. Isola et al. (2003), we have
$a_{1}=-\frac{\beta}{m_{P}},$ (50)
[…]
Figure 4: The Feynman diagrams of the $\pi^{+}$-$\pi^{+}$, $K^{+}$-$K^{+}$ and $D^{+}$-$D^{+}$ interactions.
In FIG. 4, we show the diagrams corresponding to the strong interactions of $\pi^{+}\pi^{+}$, $K^{+}K^{+}$ and $D^{+}D^{+}$, which occur through the exchange of $\rho^{0}$, $\omega$ and $\phi$. The effective potentials read
$V_{\pi^{+}\pi^{+}}$ $=$ $2g_{V}^{2}Y(\Lambda,m_{\rho^{0}},r),$ (51)
$V_{K^{+}K^{+}}$ $=$ $\frac{g_{V}^{2}}{2}\left[Y(\Lambda,m_{\rho^{0}},r)+Y(\Lambda,m_{\omega},r)\right.$ (52)
$\left.+2Y(\Lambda,m_{\phi},r)\right],$
$V_{D^{+}D^{+}}$ $=$ $\frac{a_{1}^{2}g_{V}^{2}m_{D^{+}}^{2}}{4}\left[Y(\Lambda,m_{\rho^{0}},r)+Y(\Lambda,m_{\omega},r)\right],$ (53)
of which the procedure of calculation is show in the appendix. Note that we have introduce the exponential form factor $F(q^{2})=e^{(q_{0}^{2}-\vec{q}^{2})/\Lambda^{2}}$ for each vetex in the diagrams. $\Lambda$ is the cutoff which is taken as 1956.95 (1000 $\rm{MeV}$ in the natural units) in the present work. The line shapes of the effective potentials in Eq. (51)-(53) are shown in FIG. 5.
Figure 5: The strong interaction potential of $V_{\pi^{+}\pi^{+}}$, $V_{K^{+}K^{+}}$ and $V_{D^{+}D^{+}}$.
The next step is using the varitional method, where we only need to replace the potential of the two mesons for hydrogenlike molecular ion and molecule, i.e., $\frac{1}{R}$, with $\frac{1}{R}+V_{\pi^{+}\pi^{+}}$, $\frac{1}{R}+V_{K^{+}K^{+}}$ and $\frac{1}{R}+V_{D^{+}D^{+}}$, respectively. From the numerical results, we see that the strong interactions between the two mesons almost do not affect the binding energies of the hydrogenlike molecular ion and molecule. The reason is that the electron(s) in the hydrogenlike molecular ion and molecule systems do not participate in the strong interaction, and these systems are bound by the electromagnetic attractive force of the electron on the mesons. The bond lengths for these systems are $2.003$ and $1.414$, respectively, in which case the strong interaction is negligible. So it does not affect the systems.
V Summary
Exploration of material structure is extremely important in physics, which helps us understand the interaction of fundamental elements in the entire world.
In the present work, we propose three types of hydrogenlike structures by replacing the proton(s) in hydrogen atom, molecular ion and molecule with meson(s) of $\pi^{+}$, $K^{+}$ or $D^{+}$. In order to study these systems, we solve the Schrödinger equation for hydrogenlike atoms, and use the variational method for hydrogenlike molecular ion and molecule. Since in the hydrogenlike matter, the masses of the mesons are all much heavier than that of electron, just like the case of hydrogen atom, molecular ion and molecule, the structures and the binding energies are almost the same as those of hydrogen systems. For $\pi^{+}e^{-}$, $K^{+}e^{-}$ or $D^{+}e^{-}$ system, the binding energy is $E_{n}=-\frac{1}{2n^{2}}$, and the most probable radius is $r_{n}=n^{2}$. For the systems of $\pi^{+}\pi^{+}e^{-}$, $K^{+}K^{+}e^{-}$ and $D^{+}D^{+}e^{-}$, there exist bound states if the spatial wave function is symmetric, while no bound states appear if the spatial wave function is antisymmetric. The binding energy of these hydrogenlike molecular ion is $-0.587$, and the bond length is $2.003$. For $\pi^{+}\pi^{+}e^{-}e^{-}$, $K^{+}K^{+}e^{-}e^{-}$ and $D^{+}D^{+}e^{-}e^{-}$ systems, if the spin wave function of the electrons is antisymmetric, and the spatial wave function is symmetric, there exist bound states. In this case, the two electrons can get close to each other, and form electronic cloud with high density near the midpoint of the two mesons, which provides electromagnetic attraction on the two mesons. As a result, the systems combine together. The binding energy of these systems is $-1.193$, and the bond length is $1.414$. Note that if the spin wave function is symmetric while the spatial wave function is antisymmetric, there is no bound state.
The hydrogenlike matter predicted in the present work has not yet paid much attention, since it is challenging to search for it experimentally. However, we believe that these types of matter is still of much interest, and would be observed in the future with the improvement of accuracy in the high energy physical experiments.
Acknowledgments
This project is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 11965016, 12335001, 11705069 and 12247101, and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), the Natural Science Foundation of Gansu Province (No. 22JR5RA389, No.25JRRA799), and the ‘111 Center’ under Grant No. B20063.
Jun-Feng Wang and Zi-Yue Cui are the first and co-first authors, they contribute to this work equally. Cheng-Qun Pang and Zhi-Feng Sun are both the corresponding authors.
Appendix: Breit approximation, Fourier transformation, and the function of $Y(\Lambda,m,r)$
After obtaining the scattering amplitude of FIG. 4, we use the Breit approximation to get the effective potential in the momentum space
$\mathcal{V}^{AB\to CD}(\textbf{{q}})=\frac{\mathcal{M}^{AB\to CD}(\textbf{{q}})}{\sqrt{\prod_{i}2m_{i}\prod_{f}2m_{f}}}.$ (54)
$m_{i}$ and $m_{f}$ are the mass of the initial and the final states. $\mathcal{V}^{AB\to CD}(\textbf{{q}})$ and $\mathcal{M}^{AB\to CD}(\textbf{{q}})$ are the effective potential in momentum space and the scattering amplitude of $AB\to CD$ process.
The effective potentials in the momentum and coordinate space are simply connected by Fourier transformation.
$\mathcal{V}^{AB\to CD}(\textbf{{r}})=\int\frac{d^{3}\textbf{{q}}}{(2\pi)^{3}}e^{i\textbf{{q}}\cdot\textbf{{r}}}\mathcal{V}^{AB\to CD}(\textbf{{q}})F^{2}(q^{2}).$ (55)
It is worth noting that the form factor $F(q^{2})$ is introduced to suppress the contribution of high momentum. In this paper, we take the exponentially parameterized form factor, i.e.,
$F(q^{2})=e^{q^{2}/\Lambda^{2}}=e^{(q_{0}^{2}-\vec{q}^{2})/\Lambda^{2}},$ (56)
where $\Lambda$ is the cutoff.
When performing the Fourier transformation, we will meet the following integral
$Y(\Lambda,m,r)=\int\frac{d^{3}q}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}\frac{1}{\vec{q}^{2}+m^{2}-i\epsilon}e^{2(q_{0}^{2}-q^{2})/\Lambda^{2}}.$ (57)
In Ref. He et al. (2024); Sun et al. (2024), we has calculated the $Y(\Lambda,m,r)$ function, which can be simplified as
$Y(\Lambda,m,r)$ $=$ $-\frac{e^{2q_{0}^{2}/\Lambda^{2}}}{(2\pi)^{2}r}\frac{\partial}{\partial r}\Bigg{\{}e^{2m^{2}/\Lambda^{2}}\frac{\pi}{2m}\Bigg{[}e^{mr}$ (58)
$+e^{-mr}-e^{mr}\text{erf}\left(\frac{\Lambda r}{2\sqrt{2}}+\frac{\sqrt{2}m}{\Lambda}\right)$
$+e^{-mr}\text{erf}\left(\frac{\Lambda r}{2\sqrt{2}}-\frac{\sqrt{2}m}{\Lambda}\right)\Bigg{]}\Bigg{\}}.$